A given wire of resistance 10Ω has a length of 5m and a cross sectional area of 4.0×10^-3m^2, Calculate the conductivity of the wire
1
Expert's answer
2020-05-27T10:41:28-0400

Let's first find the resistivity of the wire from the formula:


"\rho =R \dfrac{A}{l},"

here, "R = 10 \Omega" is the resistance of the wire, "A = 4.0 \cdot 10^{-3} m^2" is the cross-sectional area of the wire and "l = 5 m" is the length of the wire.

Then, we get:


"\rho = 10 \Omega \cdot \dfrac{4.0 \cdot 10^{-3} m^2}{5m} = 8.0 \cdot 10^{-3} \Omega \cdot m."

Finally, we can find the conductivity of the wire from the formula:


"\sigma = \dfrac{1}{\rho} = \dfrac{1}{8.0 \cdot 10^{-3} \Omega \cdot m} = 125 \dfrac{S}{m}."

Answer:

"\sigma = 125 \dfrac{S}{m}."